OpenStudy (anonymous):

Find the derivative of the function. y=x/sqrt(x^2+1)

OpenStudy (anonymous):

use uv rule

OpenStudy (anonymous):

\[y=x/\sqrt{x ^{2}+1}\]

OpenStudy (anonymous):

well I understand part of it

OpenStudy (anonymous):

use quotient rule

OpenStudy (anonymous):

U-substitution, yo. Or whatever they call it nowadays.

OpenStudy (anonymous):

then gemeraal power rule

OpenStudy (anonymous):

general

OpenStudy (anonymous):

Oh, wait, derivative, not integral. XD

OpenStudy (anonymous):

i get to this answer

OpenStudy (anonymous):

y=x/sqrt(x^2+1) Use Quenient and Chain rule

OpenStudy (anonymous):

Alright, product rule of x*(x^2+1)^(-1/2), (x^2+1)^(-1/2)-(1/2)x(2x)(x^2+1)^(-3/2) Yeah.

OpenStudy (anonymous):

\[y'=(x ^{2}+1)^{1/2} - x ^{2}(x ^{2}+1)^{-1/2}/x ^{2}+1\]

OpenStudy (anonymous):

thqats the derivative now i have to figure out how to simplify it

OpenStudy (anonymous):

is the answer

OpenStudy (anonymous):

oh crud made a mistake (sqrt(x^2+1)) - x(2x/2sqrt((x^(2)+1)))/(sqrt(x^2+1))^(2)

OpenStudy (anonymous):

that is the answer

OpenStudy (anonymous):

because the final answer in the back of the book is \[1/\sqrt{(x ^{2}+1)^{3}}\]

OpenStudy (anonymous):

thats the answer the book gives me

OpenStudy (anonymous):

i know my first answer is correct but not simplified

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

if i understood how to simplify it I wouldnt be here

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

fair enough :) i will help or try to my battery is about to die

OpenStudy (anonymous):

( (sqrt(x^2+1)) - x(2x/2sqrt((x^(2)+1))) )/(sqrt(x^2+1))^(2) = ( (sqrt(x^2+1)) - (x/sqrt(x^(2)+1)) )/ (sqrt(x^2+1))^(2) = (sqrt(x^2+1)) / (sqrt(x^2+1))^(2) - ( (x/sqrt(x^(2)+1)) / (sqrt(x^2+1))^(2) ) = 1 / (sqrt(x^2+1))^(2) - ( (x/sqrt(x^(2)+1)) / (sqrt(x^2+1))^(2) )

OpenStudy (anonymous):

Truly, simplification and algebraic manipulation are the difficult parts of calculus; not the class' own namesake.

OpenStudy (mimi_x3):

Lol, yes it is in the quotient rule, thats why i hate it. xD

OpenStudy (anonymous):

somehow it simplifies to

OpenStudy (anonymous):

1 / (sqrt(x^2+1))^(2) - ( (x/sqrt(x^(2)+1)) / (sqrt(x^2+1))^(2) ) = 1 / (sqrt(x^2+1))^(2) - x(sqrt(x^2+1))^(2) )/sqrt(x^(2)+1) = 1 / (sqrt(x^2+1))^(2) - x(sqrt(x^2+1))/1

OpenStudy (anonymous):

1 / (sqrt(x^2+1)) - x(sqrt(x^2+1)) sorry made a mistake

OpenStudy (anonymous):

\[(x ^{2}+1)^{-3/2}(x ^{2}+1)/(x^2=1)\]

OpenStudy (anonymous):

/x^2+1) oops

OpenStudy (anonymous):

But yeah try multiplying both the top and the bottom by the conjugate: (x(sqrt(x^2+1))^(2) + 1) but as far as I would go to simplify this would be it it ( x(sqrt(x^2+1))^(2) - 1 )/(sqrt(x^2+1))

OpenStudy (anonymous):

for final answer power in denominator should be 3/2

OpenStudy (anonymous):

\[(\sqrt{x ^{2}+1} - x^2/\sqrt{x^2+1})/(x^2+1)\]

OpenStudy (anonymous):

\[(x^2+1-x^2)/\sqrt{x^2+1}/x^2+1\]

OpenStudy (anonymous):

\[1/(x^2+1)(\sqrt{x^2+1})\]

OpenStudy (anonymous):

would \[(x^2+1)(\sqrt{x^2+1})= \sqrt{(x^2+1)^3}\]

OpenStudy (anonymous):

because the final answer should be 1/sqrt(x+1)^3

OpenStudy (anonymous):

i understand that marina same thing i have there

OpenStudy (anonymous):

(x+1)^2/3 = sqrt((x+1)^2)

OpenStudy (anonymous):

^3 oops

OpenStudy (diyadiya):

you're right y\[(x^2+1) \sqrt{x^2+1}= (x^2+1)(x^2+1)^{1/2}= (x^2+1)^{3/2}\]

OpenStudy (anonymous):

i do believe \[(x^2+1)(\sqrt{x^2+1}) = \sqrt{(x^2+1)^3}\]

OpenStudy (anonymous):

ok :) then the way i did that is right

OpenStudy (diyadiya):

Right!